Skip to main content

Performance: Some Exchange Quota Errors

Last Updated:
2012-08-17 00:00:00
Event:
2012-08-17 00:00:00
Status:
Closed
Brief Description:
User Impact:
N/A
Workaround:
There is no workaround for this issue
Current Status:
N/A
Services Affected:
Full Description:
There are some Exchange accounts that were affected by incorrect mailbox size quotas. This is currently being corrected.



Timeline of Changes

Description Current Status Date Time
On Friday afternoon, we received multiple reports of the inability to send and receive email for Exchange email users who are classified as both staff or faculty and students. Staff mailboxes have an upper limit of 7 Gb. Undergraduate mailboxes have a limit of 10 Mb. The wrong mailbox size has been applied to a number of staff and faculty accounts causing the mailboxes to be disabled. \nWe are continuing to work on resolving the issue. On Friday afternoon, we received multiple reports of the inability to send and receive email for Exchange email users who are classified as both staff or faculty and students. Staff mailboxes have an upper limit of 7 Gb. Undergraduate mailboxes have a limit of 10 Mb. The wrong mailbox size has been applied to a number of staff and faculty accounts causing the mailboxes to be disabled. \nWe are continuing to work on resolving the issue. 2012-08-17 00:00:00
All faculty and staff Exchange accounts have been returned to 7 GB limits and can now send and receive email as of about 4:45 pm. Please report any problems to the service desk at 255-5500 option 1. All faculty and staff Exchange accounts have been returned to 7 GB limits and can now send and receive email as of about 4:45 pm. Please report any problems to the service desk at 255-5500 option 1. 2012-08-17 00:00:00
Mail sent to affected email accounts was returned to the senders between about 2 pm and 4:45 pm on Friday. Mail sent to affected email accounts was returned to the senders between about 2 pm and 4:45 pm on Friday. 2012-08-17 00:00:00
This was resolved at 4:45 PM and there have been no further reports of problems concerning this issue. This was resolved at 4:45 PM and there have been no further reports of problems concerning this issue. 2012-08-17 00:00:00